Principles of Scientific Computing Preface

Jonathan Goodman last revised January 2, 2003

This book grew out of a course in Scientific Computing for graduate students at New York University. It covers a common core of material that most computing practitioners should know.

Contents

Preface.

1. Introduction.

2. Sources of error.

2.1 Computer arithmetic 2.2 Taylor approximation 2.3 Iterative methods 2.4 Statistical error in Monte Carlo 2.5 Conditioning and error amplification

3. Conditioning and accuracy